# Channel Coding for Unequal Error Protection in Digital Semantic Communication

### v1

@@ Abstract
Semantic communication is an emerging paradigm that prioritizes transmitting task-relevant information over accurately delivering raw data bits. In this paper, we address an unequal error protection (UEP) problem in digital semantic communication, where bits of higher semantic importance require stronger protection. To quantify bit-level importance, we leverage bit-flip probabilities of semantic bits as target error protection levels, which are jointly learned with semantic encoder and decoder. We propose two novel channel coding frameworks aimed at minimizing the total blocklength while satisfying UEP constraints. First, we develop a bit-level UEP framework based on repetition coding, in which the repetition number for each bit is optimized to precisely meet its target bit-flip probability. Second, we introduce a block-level UEP framework utilizing modern channel codes, where semantic bits with similar target bit-flip probabilities are grouped to exploit coding gains. Within this framework, we propose a bit-grouping algorithm guided by finite blocklength capacity analysis. Simulation results conducted on image transmission tasks confirm that the proposed frameworks significantly outperform conventional approaches, yielding substantial improvements in both task performance and transmission efficiency.
Index Terms:
Semantic communication, Semantic importance, Unequal error protection, Channel coding, Repetition coding
@@ I Introduction
Channel coding has long been a key ingredient for addressing the unequal error protection (UEP) problem in traditional communication systems [19, 20, 21]. For instance, in [19], rate-compatible punctured convolutional codes (RCPC) were proposed to assign higher code rates to less important information blocks. However, the exact error probability analysis of convolutional codes is intractable, making real-time rate adaptation difficult in practice. In [20], the assignment of code rates was optimized to different source blocks for progressive transmission using dynamic programming. While the method achieves optimality, it incurs excessive computational overhead, and estimating the expected distortion of each state remains a challenge, particularly in the context of semantic communication. In [21], it was shown that certain bit positions inherently possess stronger error-correction capabilities, and formal conditions were derived to characterize this behavior. Nonetheless, exploiting this potential requires maximum-likelihood decoding, thereby limiting the use of efficient bounded-distance decoding algorithms such as the Berlekamp-Massey algorithm [22].
Only a limited number of studies have investigated the UEP problem in the context of semantic communication [23, 24]. In [23], the authors addressed UEP for multi-modal semantic communication by applying the robust verification problem (RVP), which identifies an input feature as important if small perturbations lead to large deviations at the decoder output. In [24], the bits corresponding to segmentation maps were considered less important than those of latent maps, and the inherent UEP capability of polar codes was leveraged to provide stronger protection for the latent map bits. However, all these methods offer only modality-level or source-level error protection, rather than bit-level granularity. Addressing bit-level error protection is critical in digital semantic communication, as it enables fine-grained control over channel coding and provides greater potential for improving coding efficiency and resource utilization. To the best of the authors’ knowledge, no prior work has explored the bit-level UEP problem in digital semantic communication.
@@
To fill this research gap, we present novel channel coding frameworks that address the bit-level UEP problem in digital semantic communication. In these frameworks, we extract the target error protection levels of semantic bits by leveraging their bit-flip probabilities, jointly learned with the semantic encoder and decoder. Based on the extracted levels, we formulate the bit-level UEP problem as the minimization of the total blocklength, subject to the constraint that each semantic bit’s actual bit-flip probability does not exceed its target value. To solve this problem, we propose a repetition coding-based framework that optimizes the number of repetitions assigned to each semantic bit. To further enhance coding gain, we also develop a block-level UEP framework utilizing modern channel codes such as polar codes and low-density parity-check (LDPC) codes. In this framework, we apply block coding to groups of semantic bits with similar target bit-flip probabilities. Simulation results demonstrate that the proposed frameworks significantly improve both task performance and transmission efficiency in digital semantic communication systems. The main contributions of this paper are summarized as follows:
@@
  • We formulate a novel channel coding problem for digital semantic communication that explicitly incorporates the inherent UEP nature of semantic bits. The objective is to minimize total blocklength while ensuring that each bit’s flip probability does not exceed its respective target. Leveraging the learned bit-flip probabilities of semantic bits as target protection levels is a new perspective that has not been explored in the existing literature.
  • We propose a bit-level UEP framework based on repetition coding, in which the number of repetitions assigned to each semantic bit is optimized to satisfy its target bit-flip probability. To solve this efficiently, we develop a low-complexity bisection algorithm for determining the minimum required repetition count per bit. This is the first work to apply repetition coding for enforcing bit-level semantic importance in digital semantic communication.
  • We present a block-level UEP framework in which semantic bits with similar protection levels are grouped and encoded using polar codes or LDPC codes. To support this, we introduce a bit-grouping algorithm guided by finite blocklength capacity analysis. This framework is the first to address bit grouping for UEP from a finite blocklength perspective in semantic communication systems.
@@
The remainder of the paper is organized as follows. In Sec. II, we describe the system model of a digital semantic communication scenario and introduce an end-to-end training method to extract bit-wise error protection levels. In Sec. III, we formulate a bit-level UEP problem and propose a repetition coding framework to solve this problem. In Sec. IV, we develop a block-level UEP framework that groups semantic bits with similar protection levels. In Sec. V, we provide simulation results that validate the effectiveness of the proposed UEP frameworks across various datasets. Finally, in Sec. VI, we present our conclusions and future research directions.
II System Model and Preliminary
In this section, we describe the digital semantic communication scenario considered in this work. We then introduce the end-to-end training method proposed in [18], which is employed to extract the target error protection levels of semantic bits.
II-A Digital Semantic Communication Scenario
@@
In this work, we consider a digital semantic communication system for image transmission [^1] In this work, we focus on an image transmission task for simplicity; however, the core components of the proposed framework are not limited to this setting and can be readily extended to other types of semantic sources and tasks., as illustrated in Fig. 1. Given an image data $\bm{u}\in\mathbb{R}^{U}$, the transmitter extracts a semantic feature from the input data using a semantic encoder as follows:
@@
We consider an additive white Gaussian noise (AWGN) channel to model the wireless link between the transmitter and the receiver. Under this channel model, the received signal at time slot $t$ is given by
$y_{t}=\sqrt{p_{t}}x_{t}+v_{t},\quad t\in[T],$ (4)
where $v_{t}\sim\mathcal{CN}(0,\sigma^{2})$ denotes Gaussian noise with zero mean and variance $\sigma^{2}$, and $p_{t}$ is the transmission power allocated to the $t$-th symbol.
@@
where $g_{\rm dec}$ is the channel decoder, and ${\bm{y}}=[y_{1},\cdots,y_{T}]^{\sf T}$ is a received signal vector. In our work, we assume that symbol demodulation is performed as a part of channel decoding. After reconstructing the semantic bit sequence, the receiver recovers each semantic feature through the dequantization process ${\sf Q}^{-1}$, as follows:
@@
Figure 2: An illustration of the training procedure of the end-to-end training method in [18] along with the resulting bit-flip probabilities for the MNIST dataset when $\lambda=10^{-4}$.
II-B Error Protection Level Extraction via End-to-End Training
In this work, we adopt the end-to-end training method proposed in [18] to extract the target error protection levels of the semantic bits in $\bm{b}$}. In this method, the entire digital communication process, from the transmission of $\bm{b}$ at the transmitter to the reconstruction of $\hat{\bm{b}}$ at the receiver, is encapsulated as $K$ parallel BSCs, where the $i$-th channel has a bit-flip probability $\mu_{i}$. Then, the set of bit-flip probabilities, $\{{\mu_{i}}\}_{i=1}^{K}$, are jointly optimized with the semantic encoder and decoder based on the following loss function [18]:
@@ II-B Error Protection Level Extraction via End-to-End Training
The end-to-end training method in [18] effectively learns the bit-flip probabilities of the semantic bits while accounting for the impact of each bit’s flip probability on task performance. Moreover, the semantic encoder learns a mapping that assigns important features to semantic bits associated with lower bit-flip probabilities. In other words, this framework jointly optimizes the feature-to-bit mapping and the corresponding bit-flip probability values. Fig. 2 illustrates the training procedure of this method, along with the resulting bit-flip probabilities for the MNIST dataset when $\lambda=10^{-4}$. Inspired by the ability of this method to capture bit-level importance, we adopt the learned bit-flip probabilities as target error protection levels in our framework.
Numerical Example (Validation of Unequal Error Protection Levels): To verify the validity of the unequal error protection levels extracted using the method in [18], we present a numerical example that demonstrates the impact of protecting different semantic bits on task performance. In this example, the semantic bit sequence of a test image is sorted in ascending order of the learned bit-flip probabilities. The sorted sequence is then divided into eight segments, with the first segment containing the most critical bits. To evaluate semantic robustness, we flip the bits in each segment individually while keeping the remaining segments unchanged. Fig. 3 visualizes reconstructed CIFAR-10 images and the corresponding peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) values when the $i$-th segment is flipped. As $i$ increases, the bits tend to have higher bit-flip probabilities, indicating lower semantic importance for the downstream task. Accordingly, both PSNR and SSIM improve with increasing $i$. Notably, flipping the first segment causes a substantial degradation in reconstruction quality, even resulting in a negative SSIM value, whereas flipping the eighth segment has minimal impact on the image. These results confirm that the semantic encoder learns to associate semantic importance with varying bit-flip probabilities, effectively translating them into the required levels of error protection.
@@
III Channel Coding for Bit-Level UEP
in Semantic Communication
In this section, we formulate the bit-level UEP problem for semantic communication, in which the learned bit-flip probabilities are set as target error protection levels. We then present a channel coding method to solve this problem based on repetition coding.
III-A Problem Formulation for Bit-Level UEP
@@
III-B Channel Coding for Bit-Level UEP
@@
IV Channel Coding for Block-Level UEP
in Semantic Communication
Although repetition coding offers analytical tractability for satisfying the bit-level UEP constraint, it provides no channel coding gain and exhibits poor transmission efficiency compared to modern channel codes. In semantic communication based on the method in Sec. II-B, certain semantic bits may share similar target bit-flip probabilities. This observation suggests that applying block coding to such groups of bits can yield improved coding gain without significantly violating the UEP constraint. Motivated by this insight, we propose a block-level UEP framework that enables grouping of bits with similar target bit-flip probabilities to exploit block coding gains. The overall structure of the proposed framework is illustrated in Fig. 4.
@@ IV-A Problem Formulation for Block-Level UEP
In the proposed block-level UEP framework, we partition the $K$ semantic bits into $G$ groups based on their target bit-flip probabilities. Let $\mathcal{K}_{i}$ denote the set of indices corresponding to the $i$-th group, consisting of $k_{i}$ bits, given by
@@
where $\bar{k}_{i-1}=\sum_{j=1}^{i-1}k_{j}$. We then jointly encode the semantic bits within each group using a block channel encoder, without explicitly applying the UEP principle. Instead, the encoder is designed to satisfy the most stringent protection level within the group, i.e., to support the minimum target bit-flip probability among the bits in the group. Specifically, an $(n_{i},k_{i})$ block code is applied to the semantic bits in group $\mathcal{K}_{i}$ such that the resulting bit error probability, denoted by $P_{\rm ib}(\mathcal{K}_{i},n_{i})$, satisfies the following inequality:
@@
Considering the constraints on bit grouping and coding rate, we formulate the block-level UEP problem as follows:
@@
where $\overline{\mathcal{K}}$ is the set of possible group sizes before encoding, and $\mathcal{N}(|\mathcal{K}_{i}|)$ is the set of possible group sizes after encoding given $|\mathcal{K}_{i}|$. In the above problem, the constraint (17c) makes all groups disjoint from each other, and (17d) makes the groups cover all bits. Note that the constraints on possible group sizes, $\mathcal{K}$ and $\mathcal{N}(|\mathcal{K}_{i}|)$, depend on the channel coding scheme applied.
IV-B Bit Grouping Algorithm
As seen in the block-level UEP problem (P2), proper grouping of the semantic bits is critical in maximizing the performance of the block coding approach. For example, when bits have similar target bit-flip probabilities, forming larger groups can reduce the overall blocklength and improve coding efficiency. In contrast, when the target probabilities differ significantly, using smaller groups allows for more precise adherence to individual protection requirements, thereby enhancing overall UEP performance. We formalize this intuition by using the normal approximation of the finite blocklength capacity, which approximates the rate $k/n$ of an $(n,k)$ code as

### v2

# Unequal Error Protection for Digital Semantic Communication with Channel Coding

@@ Abstract
This paper investigates unequal error protection (UEP) in digital semantic communication, where semantically important bits require substantially higher reliability than less critical ones. To characterize this heterogeneity, we introduce a novel perspective that treats learned bit-flip probabilities of semantic bits as target error protection levels, thereby directly linking semantic importance to bit-level reliability. This formulation reveals that the required protection levels of the semantic bits may differ by several orders of magnitude, making short-block coding more advantageous than conventional long-block designs. Motivated by this, we develop two UEP frameworks that minimize total blocklength under heterogeneous reliability constraints. First, we propose a bit-level UEP framework based on repetition coding, providing an analytically tractable solution that precisely meets per-bit protection requirements. Second, to improve energy and blocklength efficiency, we design a block-level UEP framework in which the semantic bits are partitioned into short blocks with similar protection levels. Guided by finite blocklength capacity analysis, we derive a closed-form threshold condition for beneficial partitioning and develop a systematic algorithm for integrating modern channel codes. Simulation results on image transmission tasks demonstrate substantial gains in both task performance and transmission efficiency compared with conventional equal-protection schemes.
@@ I Introduction
Channel coding has long been a key ingredient for addressing the unequal error protection (UEP) problem in traditional communication systems [10, 7, 29, 18]. For instance, in [10], rate-compatible punctured convolutional codes (RCPC) were proposed to assign higher code rates to less important information blocks. However, the exact error probability analysis of convolutional codes is intractable, making real-time rate adaptation difficult in practice. In [7, 29], the assignment of code rates was optimized to different source blocks for progressive transmission using dynamic programming. While the method achieves optimality, it incurs excessive computational overhead, and estimating the expected distortion of each state remains a challenge, particularly in the context of semantic communication. In [18], it was shown that certain bit positions inherently possess stronger error-correction capabilities, and formal conditions were derived to characterize this behavior. Nonetheless, exploiting this potential requires maximum-likelihood decoding, thereby limiting the use of efficient bounded-distance decoding algorithms such as the Berlekamp-Massey algorithm [13].
However, the practical application of UEP has often been hindered by two main challenges. First, as discussed in [10, 7, 29, 18], traditional UEP schemes incur excessive computational overhead or analytical intractability. Second, in conventional source data, the disparity in importance is typically not extreme enough to justify the loss of coding gain that occurs when a single long message vector is partitioned into several short message vectors [5]. As we later establish in Theorem 1, splitting a block is only beneficial when the difference in required protection levels between bit blocks exceeds a certain threshold. In many traditional scenarios, this condition is rarely met, making the coding gain of a long block more attractive than the granularity of UEP. However, this trade-off can be exploited beneficially in semantic communication scenarios, where the importance of task-relevant bits can vary by several orders of magnitude [20].
Despite this potential, only a limited number of studies have investigated the UEP problem in the context of semantic communication [11, 9, 35]. In [11], the authors addressed UEP for multi-modal semantic communication by applying the robust verification problem (RVP), which identifies an input feature as important if small perturbations lead to large deviations at the decoder output. In [9], the bits corresponding to segmentation maps were considered less important than those of latent maps, and the inherent UEP capability of polar codes was leveraged to provide stronger protection for the latent map bits. In [35], the authors utilized the attention score of each word to classify its importance into three levels, so that more important words are protected by a more robust BCH code. However, all these methods offer only modality-level, source-level or word-level error protection, rather than bit-level granularity. Addressing bit-level error protection is critical in digital semantic communication, as it enables fine-grained control over channel coding and provides greater potential for improving coding efficiency and resource utilization. To the best of the authors’ knowledge, no prior work has explored the bit-level UEP problem in digital semantic communication.
@@
To bridge this research gap, we introduce a novel perspective that treats the learned bit-flip probabilities of semantic bits as target error protection levels, thereby directly linking semantic importance to bit-level reliability. This formulation reveals that the required protection levels of semantic bits may differ by several orders of magnitude, making short-block coding more advantageous than conventional long-block designs. Inspired by this observation, we develop two UEP frameworks that minimize the total blocklength while ensuring that the actual bit-flip probability of each semantic bit does not exceed its corresponding target value. Simulation results demonstrate that the proposed frameworks significantly improve both task performance and transmission efficiency in digital semantic communication systems. The main contributions of this paper are summarized as follows:
@@
  • We propose a bit-level UEP framework for digital semantic communication that explicitly captures the inherent unequal error protection characteristics of semantic bits. In this framework, we formulate a bit-level UEP problem that minimizes the total blocklength while ensuring that the flip probability of each bit does not exceed its corresponding target value. The formulated problem is solved using repetition coding, where the repetition number assigned to each semantic bit is optimized to satisfy its target bit-flip probability. Leveraging the learned bit-flip probabilities of semantic bits as target protection levels provides a novel perspective that has not been explored in the existing literature.
  • We propose a block-level UEP framework which partitions semantic bits with similar protection requirements into common blocks to exploit coding gains. Specifically, guided by finite blocklength capacity analysis, we establish that partitioning semantic bits into short blocks becomes advantageous when the difference in their required protection levels exceeds a certain threshold. By deriving this threshold in closed form, we develop a block partitioning algorithm that determines the optimal blocks for semantic bits according to their target protection levels. We further introduce a systematic method for integrating modern channel codes, such as polar and low-density parity-check (LDPC) codes, into the proposed partitioning framework. To the best of our knowledge, this is the first work to address UEP-aware block partitioning from a finite blocklength perspective in digital semantic communication systems.
@@
The remainder of the paper is organized as follows. In Sec. II, we describe the system model of a digital semantic communication scenario. In Sec. III, we formulate a bit-level UEP problem and propose a repetition coding framework to solve this problem. In Sec. IV, we develop a block-level UEP framework that partitions semantic bits into short blocks with similar protection levels. In Sec. V, we provide simulation results that validate the effectiveness of the proposed UEP frameworks across various datasets. Finally, in Sec. VI, we present our conclusions and future research directions.
II System Model
@@
In this work, we consider a digital semantic communication system for image transmission [^1] In this work, we focus on an image reconstruction task for simplicity; however, the core components of the proposed framework are not limited to this setting and can be readily extended to other types of semantic sources and tasks., as illustrated in Fig. 1. Given an image data $\bm{u}\in\mathbb{R}^{U}$, the transmitter extracts a semantic feature from the input data using a semantic encoder as follows:
@@
The received signal at time slot $t$ is given by
$\tilde{y}_{t}=h\sqrt{p_{t}}x_{t}+v_{t},\quad t\in[T],$ (4)
where $h\in\mathbb{C}$ is the complex-valued channel coefficient, $v_{t}\sim\mathcal{CN}(0,\sigma^{2})$ denotes the additive white Gaussian noise (AWGN) with zero mean and variance $\sigma^{2}$, and $p_{t}$ is the transmission power allocated to the $t$-th symbol. We assume that $h$ takes independent values when transmitting each semantic feature from different input data samples. Under this assumption, an equalized received signal after perfect channel equalization is given by $y_{t}=\tilde{y}_{t}/h$ for all $t$.
@@
where $g_{\rm dec}$ is the channel decoder, and ${\bm{y}}=[y_{1},\cdots,y_{T}]^{\sf T}$ is a received signal vector. In our work, we assume that symbol demodulation is performed as a part of channel decoding. From the reconstructed bits, the receiver recovers each semantic feature through the dequantization process ${\sf Q}^{-1}$, as follows:
@@
III Bit-Level UEP Framework for Digital Semantic Communication
In this section, we propose a bit-level UEP framework for semantic communication. To begin, we introduce a learning-based strategy to characterize target error protection levels for semantic bits. We then formulate the bit-level UEP problem that aims at guaranteeing these protection levels. We finally present a repetition coding method to solve the formulated problem. Furthermore, the repetition numbers determined by this method will serve as the basis for the initial block partitioning in the block-level UEP framework presented in Sec. IV.
III-A Characterizing Target Error Protection Levels via BSC Modeling
As described in Sec. II, each semantic bit $b_{i}$ undergoes channel encoding, transmission over the physical channel, and subsequent decoding at the receiver, resulting in $\hat{b}_{i}$. This end-to-end process can be equivalently represented by a channel transfer function $\mathcal{T}_{i}(\cdot)$, defined as
$\hat{b}_{i}=\mathcal{T}_{i}(b_{i})\in\{0,1\},\quad i\in[K],$ (8)
where $\mathcal{T}_{i}(\cdot)$ denotes the bit-level transfer function that transforms the transmitted bit $b_{i}$ into the reconstructed bit $\hat{b}_{i}$. The characteristics of $\mathcal{T}_{i}(\cdot)$ depend on the specific digital communication configuration, including the channel coefficients, noise power, channel coding scheme, and power allocation strategy. Therefore, during training, accurately modeling $\mathcal{T}_{i}(\cdot)$ is often infeasible, not only because of the absence of precise channel knowledge, but also due to the diversity of possible communication settings (e.g., coding and power allocation schemes).
To overcome this limitation, we adopt a simple yet effective abstraction by modeling these transfer functions as $K$ parallel BSCs as in [8, 2, 23]. In this model, each semantic bit $b_{i}$ is assumed to pass through an independent [^2] Different semantic bits may exhibit some degrees of correlation depending on the communication scenario (e.g., when block channel coding is applied). Nevertheless, we employ the independent BSC assumption solely to characterize the required bit-flip probability levels during the offline training phase. No such independence assumption is imposed during the inference phase. BSC characterized by a unique bit-flip probability $\mu_{i}$. We then interpret $\mu_{i}$ not as a mere physical error rate, but as a measure of semantic importance learned through end-to-end training. In other words, $\mu_{i}$ represents the target error protection level that the system learns to be maximally tolerable for that specific bit to perform the downstream task effectively. During training, the semantic encoder learns to map semantically important features to bits that are assigned to a low $\mu_{i}$. Therefore, we propose to utilize these jointly learned bit-flip probabilities $\{\mu_{i}\}_{i=1}^{K}$ as the target error protection levels during the inference stage.
Among various approaches for determining the set of bit-flip probabilities during training, we adopt the end-to-end training method proposed in [20]. Nevertheless, it is important to note that the design of the proposed UEP schemes is universal and can be applied with any choice of bit-flip probabilities. When employing the method in [20], the bit-flip probabilities of the BSCs are jointly optimized with the semantic encoder and decoder, allowing the encoder to learn a mapping that assigns task-relevant features to bits associated with lower bit-flip probabilities. In other words, the framework jointly learns the feature-to-bit mapping and the corresponding bit-flip probability values. Following [20], the set of the bit-flip probabilities, $\{\mu_{i}\}_{i=1}^{K}$, is jointly optimized with the semantic encoder and decoder using the following loss function:
@@ III-A Characterizing Target Error Protection Levels via BSC Modeling
Numerical Example (Numerical Evidence for Bit-Flip Probabilities as Target Protection Levels): In this example, we numerically demonstrate that the learned bit-flip probabilities serve as a reasonable choice for the target error protection levels. To this end, the semantic bit sequence of a test image is sorted in ascending order of the learned bit-flip probabilities. The sorted sequence is then divided into eight segments, with the first segment containing the bits with the lowest bit-flip probabilities. We flip the bits in each segment individually while keeping the remaining segments unchanged. Fig. 2 visualizes the reconstructed CIFAR-10 images along with the corresponding peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) values when the $i$-th segment is flipped. As $i$ increases, the bits tend to have higher bit-flip probabilities, indicating lower semantic importance for the downstream task. Consequently, both PSNR and SSIM improve with increasing $i$. Notably, flipping the first segment causes a severe degradation in reconstruction quality, even resulting in a negative SSIM value, whereas flipping the eighth segment has minimal impact on the reconstructed image. These results suggest that bits with lower learned bit-flip probabilities require stronger levels of error protection.
@@
III-B Problem Formulation for Bit-Level UEP
@@
III-C Simple Solution: Repetition-Coding-Based Approach
@@
IV Block-Level UEP Framework for Digital Semantic Communication
Although repetition coding offers analytical tractability for satisfying the bit-level UEP constraint, it provides no channel coding gain and exhibits poor transmission efficiency compared to modern channel codes. In semantic communication based on the method in Sec. III-A, certain semantic bits may share similar target bit-flip probabilities. This observation suggests that applying block coding to such blocks of bits can yield improved coding gain without significantly violating the UEP constraint. Motivated by this insight, we propose a block-level UEP framework that enables partitioning of bits with similar target bit-flip probabilities based on the finite blocklength capacity analysis. The overall structure of the proposed framework is illustrated in Fig. 3.
@@ IV-A Problem Formulation for Block-Level UEP
In the proposed block-level UEP framework, we partition the $K$ semantic bits into $G$ blocks based on their target bit-flip probabilities. Let $\mathcal{K}_{i}$ denote the set of indices corresponding to the $i$-th block, consisting of $k_{i}$ bits, given by
@@
where $\bar{k}_{i-1}=\sum_{j=1}^{i-1}k_{j}$. We then jointly encode the semantic bits within each block using a block channel encoder, without explicitly applying the UEP principle. Instead, the encoder is designed to satisfy the most stringent protection level within the block, i.e., to support the minimum target bit-flip probability among the bits in the block. Specifically, an $(n_{i},k_{i})$ block code is applied to the semantic bits in block $\mathcal{K}_{i}$ such that the resulting bit error probability, denoted by $P_{\rm ib}(\mathcal{K}_{i},n_{i})$, satisfies the following inequality:
@@
Considering the constraints on block partitioning and coding rate, we formulate the block-level UEP problem as follows:
@@
where $\overline{\mathcal{K}}$ is the set of possible block sizes before encoding, and $\mathcal{N}(|\mathcal{K}_{i}|)$ is the set of possible block sizes after encoding given $|\mathcal{K}_{i}|$. In the above problem, the constraint (18c) makes all blocks disjoint from each other, and (18d) makes the blocks cover all bits. Note that the constraints on possible block sizes, $\mathcal{K}$ and $\mathcal{N}(|\mathcal{K}_{i}|)$, depend on the channel coding scheme applied.
IV-B Theoretical Foundation for Block Partitioning
As seen in the block-level UEP problem (P2), proper partitioning of the semantic bits is critical in maximizing the performance of the block coding approach. For example, when bits have similar target bit-flip probabilities, forming long blocks can reduce the overall blocklength and improve coding efficiency. In contrast, when the target probabilities differ significantly, using short blocks allows for more precise adherence to individual protection requirements, thereby enhancing overall UEP performance. We formalize this intuition by using the normal approximation of the finite blocklength capacity, which approximates the rate $k/n$ of an $(n,k)$ code as